2a. Blood drawn from a healthy adult male has 24.0 mM НСО, and a PCO2 of 40.0 mm Hg. Calculate the pH of this sample. Use 0.0300 mmoVLemm Hg as Bunsens solubility coefficient for CO2 and 6.10 for pK, of carbonic acid. 2b. Blood drawn from a patient with chronic pulmonary obstruction has the following parameters 30.0 mM HCO,, and a PCO2 of 60.0 mm Hg. Calculate the pH of this sample. Use the same constants as in question 2a. 2c. One physiological response to correct the condition in question 2b is to increase plasma HCO, by reducing renal bicarbonate excretion. Assuming that the patients PCO2 remains 60.0 mm Hg. to what value would HC03. have to increase in order for the patients blood pH to be the same as the healthy individual in question 2a. Use the same constants as in question 2a.
Answer #1

2a, PH = 6.1+log (Hco2-)/.03Xpco2

6.1+log 24/.03X 40

6.1 + log 24/1.2

PH= 7.39

2b, PH = 6.1 + log ( Hco2-)/.03 X pco2

PH = 6.1 + log 30/1.8

PH = 7.31